OpenStudy (carlyleukhardt):

The temperature of a gas decreases 0.045°C every minute for 8 min. What is the change in the temperature of the gas? Use a positive or negative number to express the change in the temperature of the gas.

imqwerty (imqwerty):

let the initial temperature be T temperature drop in 1 min=0.045 temperature drop in 8mins=(0.045)8 new temperature = T-(0.045)8 change in temperature=final temp - initial temp =T-(0.045)8-T =-(0.045)8 = -0.36 negative sign indicates that the temperature reduced :)

OpenStudy (carlyleukhardt):

Omg qwerty thank you so much <33

imqwerty (imqwerty):

haha np :D